# Cascade switching current detectors based on arrays of Josephson junctions

### v2

@@ Abstract
Cascade multiplication is a common technique to enhance the sensitivity of photon detectors. In this study, we demonstrate novel cascade-amplified superconducting detectors utilizing arrays of Josephson junctions. The mutual coupling between junctions induces avalanche-like switching of multiple junctions upon photon absorption, leading to cascade amplification of the readout voltage. We present two prototypes featuring either low-$T_{c}$ linear $\mathrm{Nb/Nb_{x}Si_{1-x}/Nb}$ arrays, or high-$T_{c}$ stacked intrinsic Josephson junctions. Both devices exhibit clear antenna effects in microwave directivities, indicating good impedance matching and absorption efficiency. The combination of high absorption efficiency and large cascade amplification has the potential to produce broadband THz sensors with sensitivity exceeding $10^{13}\ \mathrm{V/W}$.
@@ Introduction
THz detectors find diverse applications, including security, environmental monitoring, medical imaging, chemical analysis, and fundamental research [1, 2, 3, 4]. Cooling to cryogenic temperatures improves detector sensitivity by reducing thermal noise and background radiation. Ultra-sensitive cryogenic detectors based on semiconductors [5], graphene [6], and superconductors [7] have been demonstrated. Superconductors are particularly well suited for THz detection because their characteristic energies (Josephson energy, superconducting gap, etc.) are naturally placed in the THz range and could be easily tuned for achieving optimal operation [8].
@@
A Josephson junction (JJ) can act as a sensitive switching current detector (SCD) with a potential for single-photon resolution in a broad microwave (MW) to THz range [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 8]. When JJ is biased just below the critical current, $I\lesssim I_{c}$, a small-amplitude electromagnetic signal triggers its switching from the superconducting to the resistive state. The responsivity of SCD can be very high, limited only by quantum fluctuations [8, 24]. The upper frequency depends on the characteristic voltage, $V_{c}=I_{c}R_{n}$, and ranges from sub-THz for low-$T_{c}$ [12], to THz for high-$T_{c}$ [21, 25, 22, 19] JJs. The highest $V_{c}\gtrsim 30\ \mathrm{mV}$ [26, 27] is achieved in intrinsic Josephson junctions (IJJs), naturally occurring in Bi2Sr2CaCu2O8+δ (Bi-2212) cuprates [28]. Operation of IJJs above 10 THz has been demonstrated [19, 25]. The atomic-scale of IJJs enables strong mutual coupling, advantageous for the creation of coherent THz sources [29, 30, 31, 32, 33, 34, 35, 19, 22].
Cascade multiplication is widely employed for enhancing detector sensitivity, e.g. in photo-multipliers and avalanche photodiodes [4]. For Josephson detectors, this could be achieved using arrays of mutually coupled JJs. The coupling leads to a current-locking phenomenon [36], where the switching of one JJ drags several neighbors into the resistive state, thereby multiplying the readout voltage [38, 36, 37]. A good detector should also have a high absorption efficiency, defined as the ratio of absorbed to incoming THz power. This necessitates impedance matching with free space, which is equally important for both THz sources and sensors [41, 39, 40, 35].
@@
In this work we demonstrate novel types of cascade-amplified SCDs featuring arrays of coupled JJs. Collective, avalanche-like switching of many JJs leads to cascade amplification of the readout voltage, boosting the SCD sensitivity. We present two prototypes based on either linear arrays of low-$T_{c}$ Nb/NbxSi1-x/Nb JJs or stacked high-$T_{c}$ IJJs in Bi-2212 whisker-type crystals. We investigate both MW and THz responses and observe non-trivial MW directivities, manifesting antenna effects associated with the sample geometry. This suggests good impedance matching and absorption efficiencies for both types of devices. We argue that the combination of high-$T_{c}$ superconductivity with high absorption efficiency and large cascade gain can yield broadband THz sensors with outstanding sensitivities.
Results
JJ arrays can be advantageously used for photon detection. It has been shown that utilization of arrays can help in impedance matching and broaden the detector range [42, 43]. However, earlier studies were focusing on heterodyne mixers, operating at the quasiparticle branch of the current-voltage ($I$-$V$) characteristics. In contrast, here we consider SCD, where the signal is generated upon switching out of the superconducting state [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 8].
Figure 1: Cascade SCD based on a linear Nb/NbxSi1-x/Nb array. (a) SEM image (left, false color) and layout (right) of the array. The studied line is marked in red. (b) Ensemble of the $I$-$V$’s without MW irradiation, obtained at slightly different bias currents. (c) The $I$-$V$’s with a fixed bias amplitude for different MW powers at $f\simeq 74.5\ \mathrm{GHz}$. Switching of multiple JJs leads to cascade amplification of the readout voltage. (d) The lock-in resistance of the array, $R$, (blue symbols, left axis) and the number of active JJs, $n$, (red line, right axis) as a function of MW power. (e) $R$ and $n$ as a function of the polarization angle, $\Theta$, for $P=\text{const}$. (f) Angular diagram of detector sensitivity. The orange line marks the orientation of the array. A profound off-axis directivity manifest the traveling-wave antenna effect.
@@ Results
Our goal is to achieve cascade amplification of the readout voltage. Individual JJs inherently exhibit variations. Without interjunction coupling, the array $I$-$V$ is simply the sum of independent JJ $I$-$V$’s. However, when the coupling is present, a current-locking phenomenon may occur [36], causing all (or several) JJs to switch simultaneously. Such collective switching leads to a cascade multiplication of the readout voltage, proportional to the number of current-locked JJs [38, 37, 36]. The efficiency of cascading depends on the uniformity of JJs and the strength of coupling. Therefore, arrays with many strongly interacting JJs are required to achieve high gain.
@@
A linearly polarized MW signal at $f\simeq 74.5\ \mathrm{GHz}$ is delivered to the sample in a quasi-optical manner. MW attenuation and polarization angle are adjusted using grid polarizers. The incoming MW power, $P$, is monitored by a Golay cell. Experimental details can be found in the Supplementary. All presented measurements are performed at zero magnetic field and $T\simeq$ 3.3 K.
@@
Fig. 1 (a) shows a scanning electron microscope (SEM) image and a layout of a $\mathrm{Nb/Nb_{x}Si_{1-x}/Nb}$ array. We study a linear section (marked red) with $N=128$ JJs. Details about fabrication and characterization of the arrays can be found in Refs. [44, 45, 46, 47]. Fig. 1 (b) shows the $I$-$V$s measured upon minor variation of the bias current amplitude, $I_{b}$, in the absence of MW. Multiple branches correspond to a different number, $n$, of active JJs. Interjunction coupling in these arrays is mediated by surface plasmons [46]. Such traveling waves are manifested by the appearance of collective resonant steps in the $I$-$V$’s [47]. The largest step is clearly visible in the rightmost (black) $I$-$V$ in Fig. 1 (b) with all $N=128$ active JJs.
Figure 2: Microwave characteristics of a Bi-2212 mesa. (a) A SEM image of a whisker-based device. (b) Sketch of a mesa formed at the intersections of the whisker with a top Au-electrode. It has a turnstile antenna geometry. (c) Two $I$-$V$s of a mesa with $N=250$ IJJs, in the absence of MW. (d) Ensemble of the $I$-$V$s for different MW powers at $I_{b}=38.1~{}\mu\mathrm{A}$. (e) The mesa resistance (symbols) and the number of active IJJs, $n$, (lines) versus the MW power, for $I_{b}=25.3$ and $38.1~{}\upmu\mathrm{A}$. (f) The directivity diagram. The blue line indicates the orientation of the whisker. The yellow background shows the large-scale electrode geometry. The four-fold directivity is consistent with the turnstile-antenna geometry of the device.
@@
Fig. 1 (c) shows a set of the $I$-$V$’s measured at a fixed $I_{b}$ for different MW powers. With increasing $P$, the switching current, $I_{s}$, decreases and progressively more JJs switch into the resistive state, leading to increasing of the readout voltage. Blue symbols and the red line in Fig. 1 (d) represent the lock-in resistance of the array, $R$, and the number of active JJs, $n$, respectively. It can be seen that the rapid increase of $R$ at low $P$ is proportional to $n$, illustrating the cascade amplification phenomenon. At high power, when all $N=128$ JJs are in the resistive state, the responsivity $dR/dP$ is greatly reduced. In this case the response is solely due to the decrease of $I_{s}(P)$, qualitatively similar to a single-JJ SCD [19, 8]. Thus the large difference in low and high- power responsivities highlights the advantage of array-based cascade SCD compared to a conventional single-JJ SCD.
A sensitive detector should efficiently absorb the incoming power. The optimal absorption (and emission) efficiency ($50\ \%$) is achieved at the impedance matching condition [40, 41, 39]. A JJ alone cannot effectively absorb radiation because its size is much smaller than the wavelength of radiation in free space, $\lambda_{0}$, [39]. Therefore, mm-long electrodes play a crucial role as impedance matching antennas [35]. The presence or absence of geometry-specific antenna effects provides a clue about the effectiveness of impedance matching and MW absorption.
@@
The active part of Nb-array is about $2\ \mathrm{mm}$ long ($\sim\lambda_{0}/2$ at 74.5 GHz). Additionally, there are long passive electrodes (without JJ) with a total length $\sim 1\ \mathrm{cm}$. These long electrodes may act as non-resonant traveling wave antennas, which are characterized by asymmetric of-axis directivities [45, 46]. Fig.1 (e) shows the array resistance (blue symbols) and the number of active JJs (red line) as a function of the polarization angle, $\Theta$, at a constant $P$. $\Theta=0$ corresponds to the MW electric field parallel to the array line. Fig. 1 (f) displays the corresponding sensitivity diagram, $\Delta R/V_{G}(\Theta)$, where $\Delta R$ is the resistance increase with respect to the turned-off MW, and $V_{G}\propto P$ is the Golay cell voltage (constant in this experiment). The diagram exhibits two lobes with a large directivity at $\Theta\simeq 50^{\circ}$ with respect to the array. Such off-axis directivity indicates traveling-wave antenna operation of long electrodes, enabling good impedance matching and absorption efficiency.
Figure 3: Detection of THz radiation from a Bi-2212 mesa. (a) The $I$-$V$ of the generator mesa for upward (blue) and downward (magenta) bias sweeps. Top axis indicates the anticipated Josephson frequency. (b) and (c) Low-bias $I$-$V$’s of the detector mesa with (c) representing the $I$-$V$ of the weak surface junction. (d) Measured detector responses versus $P_{gen}=I_{gen}V_{gen}$. The olive line is obtained with $I_{b}=25~{}\mu$A, activating only a single surface IJJ. The magenta curve is measured with $I_{b}=370~{}\mu$A, activating many IJJs. (e) A set of detector $I$-$V$’s with $I_{b}=370~{}\mu$A at different bias points of the generator, marked by the same-color circles in (a) and (d). It can be seen that the cascade amplification factor, $n$, is playing the dominant role in boosting the detector sensitivity.
@@
Figure 2 (a) shows a SEM image of one of the studied high-$T_{c}$ devices. They are based on Bi-2212 whiskers - needle like single crystals [22]. Each device contains several mesas, representing stacks of IJJs. Mesas are formed at the intersection of the whisker with top gold electrodes, as sketched in Fig. 2 (b). To reduce the $I_{c}$ and increase the sensitivity, some mesas were trimmed using focused ion beam (FIB) etching [48, 19]. Details about sample fabrication and characterization can be found in Refs. [48, 19, 22, 35] and the Supplementary.
@@
Fig. 2 (c) displays two $I$-$V$’s of a mesa ($\sim 5\times 5$ $\upmu\mathrm{m}^{2}$), containing $N\simeq 250$ IJJs, without MW. The blue curve, with $n=117$ active IJJs, is measured with $I_{b}$ slightly above the mean switching current, $I_{s}^{*}\sim 50~{}\upmu\mathrm{A}$. A multi-branch structure ends at $V\simeq 5$ V when all $N\simeq 250$ IJJ’s switch into the resistive state (red curve). The large readout voltage is caused by the large $V_{c}\simeq 20\ \mathrm{mV}$ per IJJ. Switching of the first and the last IJJ occurs within an interval, $\sigma\sim 5\ \upmu\mathrm{A}$   $\sim 0.1~{}I_{s}^{*}$, representing the width of the switching current distribution [49, 8]. By varying $I_{b}$ in this narrow range, we can continuously tune $n$ from 1 to 250. The responsivity of SCD is determined by the effective differential resistance, $R_{d}=dV/dI$, in the switching region [9]. For this mesa it is very large, $R_{d}=NV_{c}/\sigma\sim 1\ \text{M}\Omega$, as a consequence of the cascading.

### v3

@@ Abstract
Cascade multiplication is a widely used approach to enhance the sensitivity of photon detectors. Although vacuum tube and semiconductor photomultipliers achieve high signal gain and enable single-photon detection in the optical range, their performance is constrained at lower frequencies by a significant work function (1–10 eV). Superconducting detectors overcome this limitation, enabling operation in the terahertz (THz) and microwave (MW) frequency ranges. In this study, we introduce novel cascade-amplified superconducting detectors based on arrays of Josephson junctions. The interjunction coupling induces avalanche-like switching of multiple junctions upon photon absorption, resulting in cascade amplification of the readout voltage. We present two prototypes featuring either low-$T_{c}$ linear $\mathrm{Nb/Nb_{x}Si_{1-x}/Nb}$ arrays or high-$T_{c}$ stacked intrinsic Josephson junctions. Both MW and THz responses are analyzed and the advantage of the cascade detector compared to a conventional single-junction detector is demonstrated. Our findings suggest that array-based cascade superconducting detectors can be used to build broadband MW-to-THz sensors, with sensitivities exceeding $10^{13}\ \mathrm{V/W}$.
@@ Introduction
Detectors in the microwave (MW) and terahertz (THz) ranges find diverse applications, including security, environmental monitoring, medical imaging, chemical analysis, and fundamental research [1, 2, 3]. Detectors with single-photon and photon-counting resolution are needed in quantum optics and electronics [3, 4, 5, 6]. However, despite intense research in this area for the past two decades [1, 2, 3, 7, 8, 10, 4, 5, 9, 6, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] MW and THz single-photon detectors (SPDs) are not yet commercially available due to many technical challenges and materials limitations.
The main challenge is associated with small photon energies [0.41 - 41] meV for the frequency range [0.1-10] THz. This puts strong constraints on detector materials. Conventional vacuum tube and semiconductor photodetectors [3] are not operational in this range due to the large work function and band gap (typically 1-10 eV). Therefore, low-gap materials such as superconductors [5, 7, 8, 10, 4, 11, 9, 12, 6, 13, 14, 15, 16], half-metals [17, 12, 18, 19], or gap-engineered quantum dots [20, 21] are required. Furthermore, in situ gap tunability is needed for optimizing detector operation and adjusting dynamic range [15]. This can be achieved through electrostatic gating [5, 17, 11, 12, 18, 19, 20, 21], magnetic field [4, 20, 15, 6], bias [10, 15, 9, 6, 22, 23], temperature [7, 8, 10, 9, 16, 15, 22, 23], etc. A sensitive detector should have a high absorption efficiency, $\chi$, necessitating the implementation of a pickup antenna for impedance matching with free space [24, 25].
The low photon energy also sets severe demands on detector characteristics. For MW SPDs, the noise-equivalent power (NEP) should be in the zW/Hz1/2 range [20, 21, 11, 12, 16] (0.66 zW corresponds to absorption of one 1 THz photon per second). So sensitive devices cannot operate at room temperature due to large black-body radiation with a power density of about $46$ mW/cm2 and a peak frequency 17.6 THz at $T=300$ K. Therefore, cooling to cryogenic temperatures is commonly required, enabling the utilization of superconductors. Superconductivity is beneficial for building ultrasensitive detectors due to the lack of Jonson-Nyquist voltage noise, $S_{V}=[4k_{b}TR]^{1/2}$, in electrodes. Many of existing superconducting detectors, most noticeably those based on qubits [5, 4, 11, 12, 6], approach the quantum limit of sensitivity set by the Heisenberg uncertainty relation. The base element of qubits, exhibiting a quantum-mechanical behavior [26], is a Josephson junction (JJ).
@@
A current-biased JJ can act as a sensitive switching current detector (SCD) with potential for single-photon resolution in a broad MW to THz range [34, 27, 22, 28, 23, 29, 30, 31, 32, 33, 15]. The responsivity of SCD can be very high, limited only by quantum fluctuations [15, 34]. The upper frequency depends on the characteristic voltage, $V_{c}=I_{c}R_{n}$, and ranges from sub-THz for low-$T_{c}$, to THz for high-$T_{c}$ [31, 35, 32, 29] JJs. The highest $V_{c}\gtrsim 30\ \mathrm{mV}$ [36, 37] is achieved in intrinsic Josephson junctions (IJJ), which naturally occur in Bi2Sr2CaCu2O8+δ (Bi-2212) cuprates [38]. The operation of IJJs above 10 THz has been demonstrated [29, 35]. The atomic scale of IJJs enables strong mutual coupling, which is advantageous for the creation of coherent THz electronics [39, 40, 41, 42, 43, 44, 29, 32]. Mutual coupling between JJs leads to a current-locking phenomenon in JJ arrays [45], where the switching of one JJ drags several neighbors into the resistive state, thereby cascade-multiplying the readout voltage [47, 45, 46].
Figure 1: The concept of cascade amplification. (a) Clarification of NEP reduction by cascade-amplification of the detector sensitivity. Olive and red lines indicate voltage responses of a single-stage and cascade-amplified detectors as a function of the incoming power. The gray shaded region indicates the readout voltage noise. (a) Energy-phase diagram of a current-biased JJ array. Switching (escape) of a single junction (red circle) leads to avalanche-like dragging of other JJs due to the mutual inter-junction coupling (green lines). (c) The $I$-$V$ characteristics of an array with $N=1000$ Nb/NbxSi1-x/Nb JJs at $T\simeq 2.5$ K. It can be seen that all JJs switch simultaneously from the superconducting to the resistive state, leading to a large readout voltage, $V\simeq 0.45$ V. The inset shows the meander-like array layout.
@@
In this work, we present a new concept of cascade-amplified SCDs based on arrays of coupled JJs. Collective, avalanche-like switching of many JJs leads to cascade amplification of the readout voltage, boosting the SCD sensitivity. We present two prototypes based on either linear arrays of low-$T_{c}$ Nb/NbxSi1-x/Nb JJs, or stacked high-$T_{c}$ IJJs in Bi-2212 whiskers. We investigate both MW and THz responses and compare single-junction and cascade SCD operation at the same device. MW absorption efficiency is analyzed by studying polarization loss diagrams, revealing antenna effects associated with device geometry. We argue that the combination of high-$T_{c}$ superconductivity and large cascade gain can yield broadband THz sensors with outstanding sensitivities.
Operation principle
@@
Cascade multiplication is widely employed to improve detector sensitivity, e.g., in photomultipliers and avalanche photodiodes [3]. Figure 1 (a) clarifies the principle of a cascade-amplified detector. The green line shows the voltage response of a single-stage detector, as a function of the incoming THz power, $P_{THz}$. Above a threshold, $P_{THz}>P_{th}$, it grows linearly with the slope determined by the sensitivity, $S=V/P_{THz}$ (V/W). NEP corresponds to $P_{THz}$ at which the response exceeds the noise floor. The sensitivity of the cascade detector is increased in proportion to the number of stages, $S_{n}=nS_{1}$. This will reduce $NEP_{n}$, provided, the noise floor is not multiplied correspondingly. Our goal is to achieve cascade amplification of the readout voltage. Individual JJs inherently exhibit variations. Without interjunction coupling, the array $I$-$V$ is simply the sum of $I$-$V$’s of individual JJs and the response in the low-power limit is determined by the weakest (single) JJ. However, when the coupling is present, a current-locking phenomenon can occur [45], causing all (or several) JJs to switch simultaneously, as sketched in Fig. 1 (b). Collective switching leads to a cascade multiplication of the readout voltage, proportional to the number of current-locked JJs [47, 46, 45]. The efficiency of cascading depends on the uniformity of JJs and the strength of coupling. Therefore, arrays with many strongly interacting JJs are required. Fig. 1 (c) shows the $I$-$V$ of an array containing $N=1000$ Nb/NbxSi1-x/Nb JJs [55]. It demonstrates a nearly ideal current locking with all JJs switch simultaneously, leading to a large readout voltage, $V\simeq 0.45$ V.
Figure 2: Operation of an ordinary single-junction SCD (a) Calculated switching probabilities as a function of bias current for a JJ with $I_{c0}=50~{}\mu$A, $Q_{0}=100$ and $T=1$ K. The blue line - without radiation and the red line with induced THz current of $I_{THz}=0.005~{}I_{c0}$. (b) Corresponding probability densities, representing switching current histograms. (c) Calculated voltage response (red line, left axis) and sensitivity (olive line, right axis) versus incoming THz power for $I_{b}=0.972~{}I_{c0}$, as indicated by vertical arrows in (b). Based on the data from Ref. [15] for the non-resonant case, assuming $V_{c}=20$ mV, relevant for a single Bi-2212 IJJ.
Single-junction SCD
Detailed description of a single-junction SCD can be found in Ref. [15]. The JJ dynamics is equivalent to motion of a particle in a tilted washboard potential, see Fig. 1 (b). The corrugation is determined by the Josephson energy $E_{J0}=(\Phi_{0}/2\pi)I_{c0}$, where $\Phi_{0}$ is the flux quantum and $I_{c0}$ is the fluctuation-free critical current; the tilt is created by the bias current, $I$; and the coordinate, $\varphi$, is given by the Josephson phase difference. When the particle escapes from the potential well, the JJ switches into the resistive state. This occurs at a switching current $I_{s}<I_{c0}$. The premature switching is caused both by the induced THz current from the impacting radiation, and by thermal and quantum current fluctuations. Therefore, SCD requires a statistical readout - the price which has to be paid for the exceptional sensitivity [15]. However, this is true for any quantum-limited detector.
The blue lines in Figure 2 (a) show a switching probability, $G(I)$, versus bias current in the absence of radiation. Calculations are made for an underdamped JJ with the quality factor $Q_{0}=100$, $I_{c0}=50~{}\mu$A at $T=1$ K [15]. The switching current histogram, $I_{s}(I)$, is given by the probability density, $g=dG/dI$, as shown in Fig. 2 (b). It has the full-width at half-maximum [48],
$\sigma_{1}\simeq aI_{c0}\left[\frac{k_{B}T}{2E_{J0}}\right]^{2/3}.$ (1)
According to our simulations, the prefactor, $a\sim 1$, slightly depends on $E_{J0}/k_{B}T$ and varies from $a=0.57$ for $E_{J0}/k_{B}T>100$ to $a\simeq 1$ for $E_{J0}/k_{B}T\sim 10$.
Red lines in Figs. 2 (a) and (b) show corresponding quantities in the presence of radiation-induced current with the amplitude, $I_{THz}=0.005I_{c0}$. Calculations are done for the non-resonant case with the frequency much smaller than the Josephson plasma frequency [15]. The SCD is biased by a low-frequency alternating current with a constant amplitude $I_{b}$ close to the middle of the $G(I)$ step in the absence of radiation. The time-average voltage response $\Delta V$ is proportional to the vertical shift of $\Delta G(I_{b})$ at the bias amplitude, as indicated by the arrows in Fig. 2 (a)
$\Delta V_{1}\simeq V_{c}\Delta G(I_{b}).$ (2)
The red line in Fig. 2 (c) shows SCD response, $V(P_{THz})$, at $I_{b}=0.972~{}I_{c0}$, marked in Fig. 2 (a), for $V_{c}=20$ mV, relevant for Bi-2212 IJJ. The response grows approximately linearly at small $P_{THz}$ and saturates when $I_{THz}$ exceeds $2\sigma_{1}$. The external $P_{THz}$ is connected to the internal $I_{THz}$ as,
$P_{THz}=\frac{I_{THz}^{2}R_{THz}}{2\chi}.$ (3)
Here $\chi$ is the absorption efficiency and $R_{THz}$ is the real part of the THz impedance. Hereinafter, we consider the optimal impedance-matched case with maximum $\chi=0.5$ [24, 25], assuming $R_{THz}=100~{}\Omega$ [49].
The olive line (right axis) in Fig. 2 (c) shows the corresponding single-JJ SCD sensitivity, $S_{1}$. It can be estimated using Eq. (19) from Ref. [15]. Taking into account that $\sigma_{1}\simeq 1/g_{max}$, where $g_{max}$ is the peak value of the histogram, we can write,
$S_{1}\sim 2\chi V_{c}R_{THz}\frac{e^{2}}{(\hbar\omega)^{2}}\frac{I_{c0}}{\sigma_{1}}.$ (4)
For the chosen parameters, Eqs. (4) and (1) yield $S_{1}\simeq 10^{9}$ (V/W), consistent with Fig. 2 (c). However, the ultimate SCD sensitivity can be much higher, up to $S_{1}=5\times 10^{12}$ (V/W) [15]. From Eq. (4) it is seen that the SCD sensitivity is limited only by the finite width of the switching current histogram. According to Eq.(1), $\sigma_{1}$ reduces with decreasing $T$ due to the reduction of thermal fluctuations. However, it saturates below the crossover temperature to macroscopic quantum tunneling, $T^{*}\sim 10$ mK [26]. Therefore, SCD represents probably the simplest quantum-limited detector.
Multi-junction cascade SCD
It has been known that JJ arrays can be advantageously used for photon detection: arrays can help with impedance matching and broaden the dynamic range [50, 51]. However, earlier studies were focusing on heterodyne mixers, operating at the quasiparticle branch of the $I$-$V$. In contrast, here we consider SCD, where the signal is generated upon switching out of the superconducting state [22, 28, 23, 30, 15].
The cascade amplification factor of the multi-junction SCD is given by the number of current-locked JJs, $n$,
$V_{n}\simeq nV_{1},~{}~{}~{}~{}(1\leq n\leq N).$ (5)
Ideally, all $N$ junctions switch simultaneously, as shown in Fig. 1 (c). However, due to various types of nonuniformity this is not always the case. Therefore, the performance of array-based SCD depends on the distribution $n(P_{Thz})$ with the most important parameters being the starting number of JJs, $n(0)$, at $P_{THz}\rightarrow 0$ and the width of the distribution, $\sigma_{n}$.
Figure 3: Calculation of sensitivity for a cascade SCD. (a) A single junction response, characterized by the power-width $\sigma_{1}$. (b) Three examples of distribution functions of the number of active JJs versus $P_{THz}$ for an array with $N=200$ JJs. The leftmost (magenta) corresponds to a fairly uniform array with a large initial number $n(0)=50$ and a narrow power-width $\sigma_{n}=\sigma_{1}/4$. The middle (wine), to $n(0)=25$ and $\sigma_{n}=\sigma_{1}$. The rightmost (red), to the least uniform array with $n(0)=10$ and $\sigma_{n}=2\sigma_{1}$. Panels (c) and (d) show array responses, calculated from Eq. (5), (c) in the full scale and (d) at small power.
Figure 3 illustrates multi-junction SCD operation. Fig. 3 (a) shows the mean single-JJ response $V_{1}(P_{THz})$, characterized by the width $\sigma_{1}$, c.f. with the red line in Fig. 2 (c). Fig. 3 (b) represents three examples of possible $n(P_{THz})$ for arrays with $N=200$ JJs. The leftmost (magenta) shows the case of a fairly uniform array with $n(0)=50$ and with rapid activation of all JJs, $\sigma_{n}=1/4~{}\sigma_{1}$; the middle (wine) to $n(0)=25$ and $\sigma_{n}=\sigma_{1}$; and the rightmost (red) to the least homogeneous array with small $n(0)=10$ and broad $\sigma_{n}=2~{}\sigma_{1}$. Figs. 3 (c) and (d) represent corresponding array responses calculated via Eq. (5), (c) in the full power range and (d) at low $P_{THz}$. It can be seen that larger $\sigma_{n}$ broadens the dynamic range of the detector, while the sensitivity at low power is determined solely by $n(0)$,
$S_{n}(0)=S_{1}n(0).$ (6)
Additional clarifications about SCD operation, lock-in readout, statistical uncertainty and noise measurement are provided in Methods and Supplementary information.
Experimental results
Below we analyze MW and THz responses of low-$T_{c}$ Nb/NbxSi1-x/Nb JJ arrays and high-$T_{c}$ mesa structures on Bi-2212 whiskers. Fabrication and physical properties of Nb arrays were described in Refs. [52, 53, 54, 55]. Details about Bi-2212 device fabrication and characterization can be found in Refs. [56, 29, 32] and the Supplementary. All presented measurements are performed at zero magnetic field and $T\simeq$ 3.3 K. Experimental details and additional data can be found in Methods and the Supplementary.
@@
A linearly polarized MW signal at $f\simeq 74.5\ \mathrm{GHz}$ is delivered to the sample in a quasi-optical manner. MW attenuation, $\mu$, and polarization angle, $\Theta$, are adjusted using grid polarizers. The incoming MW power, $P$, is monitored by a Golay cell.
@@
We study a linear Nb array containing $N=128$ Nb/NbxSi1-x/Nb JJs. Fig. 4 (a) shows the $I$-$V$s measured upon minor variation of the bias current in the absence of MW. Multiple branches correspond to a different number, $n$, of active JJs. Interjunction coupling in these arrays is mediated by surface plasmons [54] and is manifested by the appearance of collective resonant steps in the $I$-$V$’s [55].
@@
Blue circles in Fig. 4 (b) show the array response measured at a fixed $I_{b}$ as a function of MW power. With increasing $P$, the switching current, $I_{s}$, decreases and progressively more JJs switch into the resistive state, leading to the increase of readout voltage. The red line in Fig. 4 (b) represents the number of active JJs, $n$. It can be seen that the rapid increase of $V$ at low $P$ is proportional to $n$, illustrating the cascade amplification phenomenon. At higher power, all $N=128$ JJs switch to the resistive state and the responsivity $dV/dP$ is greatly reduced. In this case the response is solely due to the decrease of $I_{s}(P)$, qualitatively similar to a single-JJ SCD [29, 15].
Figure 4: Cascade SCD based on a linear Nb/NbxSi1-x/Nb array with $N=128$ JJs. (a) Ensemble of the $I$-$V$’s without MW irradiation, obtained at slightly different bias currents. The inset shows the array layout. (b) The array response (blue symbols, left axis) and the number of active JJs, $n$, (red line, right axis) as a function of MW power. (c) $V$ and $n$ as a function of the polarization angle, $\Theta$, for $P=\text{const}$. The inset shows the polarization loss-function. The orange line marks the array direction. Profound off-axis lobes manifest the traveling-wave antenna effect.
As follows from Eq. (4), detector sensitivity is directly proportional to the absorbtion efficiency, $\chi$. A JJ alone cannot effectively absorb radiation because its size is much smaller than the wavelength in free space, $\lambda_{0}$ [25]. However, the 2 mm long active part of the array is $\sim\lambda_{0}/2$ at 74.5 GHz. There are also passive electrodes (without JJ) with the total length $\sim 1\ \mathrm{cm}$. As shown in earlier studies of MW emission from such arrays [53, 54], the long electrodes act as non-resonant traveling wave antennas, enabling impedance matching. The key signature of the traveling wave antenna is an asymmetric off-axis directivity [24, 53, 54].
@@
To analyze the absorption efficiency of the array, we measure its polarization loss factor [57]. Fig.4 (c) shows the detector voltage (blue symbols, left axis) and the number of active JJs (red line, right axis) as a function of the MW polarization angle, $\Theta$, at a constant $P$. $\Theta=0$ corresponds to the MW electric field parallel to the array line. The inset displays the corresponding polarization-loss diagram, $\Delta V/V_{G}(\Theta)$, where $\Delta V$ is the array voltage with respect to the turned-off MW, and $V_{G}\propto P$ is the Golay cell voltage (constant in this experiment). The diagram exhibits two lobes with profound maxima at $\Theta\simeq 50^{\circ}$ and $-130^{\circ}$ with respect to the array. The off-axis behavior is consistent with traveling-wave antenna operation of long electrodes, enabling good impedance matching and absorption efficiency.
Figure 5: Microwave characteristics of a Bi-2212 mesa. (a) A SEM image of a whisker-based device. (b) $I$-$V$s of a mesa with $N=250$ IJJs, measured with two different bias amplitudes without MW. (c) Red circles represent measured $T$-dependence of the width of switching current histograms for a single IJJ. The blue line is calculated from Eq. (1). The inset shows the measured switching current histogram (red) and the calculated switching probability density (blue line). (d) Ensemble of the $I$-$V$s for different MW powers, determined by the attenuation factor $\mu$, at constant $I_{b}=38.1~{}\mu\mathrm{A}$. (e) The mesa voltage (symbols) and the number of active IJJs, $n$, (lines) versus the MW power, for $I_{b}=25.3$ and $38.1~{}\upmu\mathrm{A}$. (f) The polarization-loss diagram. The blue line indicates the orientation of the whisker. The yellow background shows the large-scale electrode geometry. The four-fold symmetry is consistent with the turnstile-antenna geometry of the device, as sketched in the inset.
@@
Figure 5 (a) shows SEM image of one of the studied high-$T_{c}$ devices. They are based on Bi-2212 whiskers - needle like single crystals [32]. Each device contains several mesas, representing stacks of IJJs. Mesas are formed at the intersection of the whisker with top gold electrodes, as sketched in Fig. 5 (f). To reduce the $I_{c}$ and increase the sensitivity, some mesas were trimmed using focused ion beam (FIB) etching [56, 29].
@@
Fig. 5 (b) displays two $I$-$V$’s of a mesa ($\sim 5\times 5$ $\upmu\mathrm{m}^{2}$), containing $N\simeq 250$ IJJs, without MW. The blue curve, with $n=117$ active IJJs, is measured with $I_{b}$ slightly above the mean switching current, $I_{s}\sim 50~{}\upmu\mathrm{A}$. A multi-branch structure ends at $V\simeq 5$ V when all $N\simeq 250$ IJJ’s switch into the resistive state (red curve). The large readout voltage (5 V) is caused by the large $V_{c}\simeq 20\ \mathrm{mV}$ per IJJ. Switching of the first and the last IJJ occurs within an interval, $\sigma_{n}\sim 5\ \upmu\mathrm{A}$   $\sim 0.1~{}I_{s}$, representing the width of the switching current distribution [58, 15]. Statistical analysis of multi-junction switching can be found in the Supplementary.
Inset in Fig. 5 (c) shows a switching current histogram (red) for a single IJJ on another mesa. The solid line shows the simulated probability density at the base $T$. Red circles in Fig. 5 (c) show measured temperature dependence of the histogram width. The blue line is obtained from Eq. (2). The overall agreement is excellent, indicating that IJJs are well described by the standard formalism.